Circular No. 8357

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

      URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2004ci IN NGC 5980

Further to IAUC 8349, M. Armstrong reports his discovery, on images taken on June 16.997 UT, of an apparent supernova (mag 17.5) located at α = 15^{h}41^{m}29^{s}.83, δ = +15^{o}47′13′′.1 (equinox 2000.0), which is 10′′.1 west and 2′′.3 south of the center of NGC 5980. Nothing was visible at this location on an image taken by Armstrong on 2003 Apr. 8. Further to IAUC 8345 and 8350, S. Park and R. J. Foley also report LOSS observations of SN 2004ci via KAIT images taken on June 15.3 (mag ∼ 17.7) and 17.3 (mag ∼ 17.4), giving position end figures 29^{s}.71, 12′′.7. A KAIT image taken on June 12.4 showed nothing at this position (limiting mag ∼ 19.0).

COMETS C/2003 X5–X11 AND Y2–Y10 (SOHO) Further to IAUC 8356, K. Battams reports measurements for addi-

tional Kreutz sungrazing comets found on SOHO website C2 images by M. Boschat (C/2003 X5, X7, Y5, Y8), X.-m. Zhou (C/2003 X6, X9, X11), T. Hoffman (C/2003 X8), X. Leprette (C/2003 X10, Y2), R. Matson (2003 Y3, Y9, Y10), R. Kracht (C/2003 Y4, Y7), and J. Sachs (C/2003 Y6). C/2003 Y6 and Y7 were also visible on C3 images. The reductions by B. G. Marsden (and by R. Kisala for C/2003 Y8, Y9, Y10) and orbital elements by Marsden appear on the MPECs cited below.

Comet  2003 UT  α_{2000}  δ_{2000}  MPEC
C/2003 X5 Dec. 4.896  16^{h}46.^{m}9  −24^{o}12′  2004-L24
C/2003 X6  6.396  16 52.2  −24 20  2004-L24
C/2003 X7  7.829  16 58.6  −24 27  2004-L24
C/2003 X8  8.246  17 00.9  −24 31  2004-L24
C/2003 X9  8.621  17 02.2  −24 34  2004-L24
C/2003 X10  11.188  17 14.5  −24 41  2004-L25
C/2003 X11  13.588  17 25.7  −24 57  2004-L25
C/2003 Y2  19.621  17 53.6  −25 08  2004-L25
C/2003 Y3  19.979  17 55.0  −25 02  2004-L25
C/2003 Y4  20.729  17 58.8  −25 14  2004-L25
C/2003 Y5  22.771  18 08.5  −25 08  2004-L25
C/2003 Y6  23.571  18 14.3  −27 22  2004-L26
C/2003 Y7  24.638  18 20.0  −27 44  2004-L26
C/2003 Y8  25.288  18 19.6  −25 00  2004-L67
C/2003 Y9  25.479  18 20.7  −24 46  2004-L67
C/2003 Y10  26.064  18 23.5  −24 50  2004-L67
2004 June 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green